OpenStudy (anonymous):

help 8b^2-50. Factor do i do it 8(b^2-50) or 8(b-25)(b+25) or 8(b-5)(b+5) OR none

OpenStudy (turingtest):

none if you take out the 8 you have to factor it out of BOTH terms 50/8 is not a perfect square, so try to factor out 2 instead remember to factor it out of both terms first

OpenStudy (turingtest):

(10a+50)=10(a+5) for example

OpenStudy (amistre64):

25^2 not= 50

OpenStudy (anonymous):

ok so 2(b+25)^2?

OpenStudy (turingtest):

why did the ^2 move outside?

OpenStudy (amistre64):

\[b\sqrt{8}\pm \sqrt{50}\]

OpenStudy (anonymous):

lol I m not sure he was nt happy inside????ok so keep 2 inside?

OpenStudy (turingtest):

yeah and keep it negative and b should still have a 4 next to it

OpenStudy (anonymous):

(2b+25)(2b-25)

OpenStudy (amistre64):

25^2 not= 50 still

OpenStudy (anonymous):

no w8

OpenStudy (turingtest):

...and the 2 that you pulled out needs to stay next to the parentheses it can't just disappear

OpenStudy (amistre64):

not to mention the other fauxpauxs

OpenStudy (anonymous):

ok thats why im on here armi have patience grasshopper

OpenStudy (anonymous):

2(4b+25)(b-25)

OpenStudy (turingtest):

\[(a^2-b^2)=(a+b)(a-b)\](just to make sure it's on the screen)

OpenStudy (anonymous):

but 8 has no sq root you see where im stuck

OpenStudy (turingtest):

let's back up... 8 and 50 are not perfect squares, but they are close divide each by 2 and they will be, so we factor out a 2:\[8b^2-50=2(4b^2-25)\]that's the first step now what is a and b in your particular problem

OpenStudy (anonymous):

4 and 25?

OpenStudy (turingtest):

no, that is a^2 and b^2 so what are a and b then?

OpenStudy (anonymous):

I DONT KNOPWWWWWWWWWWWWWWWWWWWWWWWWWWWWWWW im sorry im challenged mathematically speaking

OpenStudy (anonymous):

I dont have a clue. I thought a and b was 8 and 25 what else is there to choose from?

OpenStudy (turingtest):

actually sorry my notation is confusing you perhaps the difference of squares formula is\[x^2-y^2=(x-y)(x+y)\]your formula is\[8b^2-50=2(4b^2-25)=2((2b)^2-5^2)\]so what I call x and y in the formula are 2b and 5 in your particular problem

OpenStudy (anonymous):

No its not you it's me. you have been really helpful thank you. I will apply it to the next equation. thanks<3

OpenStudy (anonymous):

It makes sense when you break it down like that<3

OpenStudy (turingtest):

welcome you know this one's not done though, you still need to apply the actual formula for difference of squares I will leave that to you

OpenStudy (anonymous):

thanks tur.